Meeting 18 January 2010

Added by Tunnell, Christopher over 13 years ago

-- minutes of previous meeting
-- review of actions
-- Matters arising from TB
-- MOM 2011 and run plan (Alain)
-- MICE magnet group
-- FAC papers and in particular the MICE MOU (Bonesini)
-- spectrometer solenoid repair plan status
--> coupling coils plans
-- conferences and papers (Palladino, Bonesini)
-- a.o.b.


Added by Tunnell, Christopher over 13 years ago

  • Alain Blondel
  • David Colling
  • Mike Zisman
  • Andy Nichols
  • John Cobb
  • Maurizio Bonesini
  • Vittorio Palladino
  • Ken Long (joined later)

Unscheduled: Blondel asks Collings how the software and computing group is going

Software is making progress and improvement in coding standards is promising. Config DB is making progress. Data mover is not working due to gLite problem.

Colling is not aware of the Online group progress since that's Jean-Sebastian Graulich's responsibility, but Yordan Karadzhov will be in Geneva to replace Graulich and they'll overlap for 4-5 months.


Blondel has sent comments on the previous meeting minutes. Blondel says that minutes have always been public, but technical challenges meant the minutes haven't been posted.

ACTION: Tunnell to find some way to print the action items


MICE website

Colling reports:

Progress made by Tunnell. We're currently working on a plan. being main face, redmine being the internal. Operations needs to be figured out as highest priority. More will be figured out later.

Blondel points out that we need to be clear (write down?) what the difference is between the usage of the elog and redmine.

ACTION UPDATE: work has been done and we need to figure out how the elog fits in

runplan 2011

Blondel sent a runplan around with the MOM schedule. We seem to have a wide range of MOMs signed up. The people side is done.

Blondel reports:

For this run period, the target and EMR are the main focuses. EMR is main priority with no rush required on the target. Exercise target in machine physics at beginning of May. Be ready for running EMR test and step 1 in July. After that, we will not take data until December.

Ken Long joins and reports that he wants one day to make sure the target is not dead.

Kaplan comments on the MOM rota that Linda Coney may be MOM too much since she's important to other stuff. Blondel says that there is a new postdoc at Imperial that could take a MOM shift. Kaplan volunteers himself as well.

ACTION: CLOSED until collaboration meeting

RF tests at cern, nothing

It's on Hold says Blondel. We need to understand the CC situation before proceeding on this says Blondel.

ACTION: CLOSED until more interest

Coupling Coils

Discussed later.

change request template in progress

Nichols talked to Jason Terrant about this today. Blondel says this is badly needed and raises the priority. People want to see a first non-perfect iteration. Derun sent MICE note, but no change request.


step 3 v 4

In progress.

solid absorbers

There needs to be a solid absorber note, so reopen. The action was closed after absorber thicknesses were shared by Tunnell.

ACTION: REOPEN and add note request

Agenda for RAL

There is a list of suggested speakers. This will get finalized soon after it gets some iteration from what Blackmore/Blondel sent. It's the job of the EB to send comments ASAP.

NEW ACTION: EB to send comments ASAP about agenda

Prepare a program of RF cavity tests

Derun will talk at VC. Long thinks this is a TB thing. The EB must have a plan, but the TB implements the plan then reports. This should be a note.

magnetic measurement plan

Blondel said that Pierrick Hanlet will get from Chris Rogers the fields and beam envelope then establish the measurement step size as a function of field variation. We also don't need to map where the muons never go.

Long says he has already suggested to sample the air in the hall to know the field map around channel so we know how the field lines go to the wall, which was an idea from Cobb prior. Cobb was told by Rogers that there may be technical problem as how to turn Cartesian into G4MICE, though Colling's conversations with Rogers haven't revealed this. It's probably a memory size issue if it isn't done cleverly.

Blondel says this is an analysis and software CM thing.

NEW ACTION on Cobb: take this to analysis meeting and find a person.

matters from TB

Courthold has now ordered a program from vector fields to understand field map. LH system UPS does not need diesel generator says ISIS, so can we can order something.

By CM, need dimensions and materials of solid absorber in hand.

One change is the way Nichols will do milestone tracking which seems to be accepted. This will be tested at Monday's MICO meeting. Nichols now has a permanent secretary for the MICO meeting. The next question is how to make MICO meeting more popular.

The proposal was made that Courthold would help with the magnets.

The SS repair plan has been presented to the MAP committee. There is concern about cost and time of review. But we can start with certain parts of the repair if allowed. Nichols says that there are quite a few tasks that are low risk associated with the magnets. Nichols wants a list that doesn't require permission. Zisman says we're been stalled by seven months because we haven't had permission to continue.

Blondel wants it to go like this: they make a proposal of things they want to do then run it through review committee so they are responsible too for it. Zisman comments that we are treating the reviewers like they are MICE management in terms of their say. Long wants it so if the TB disagrees with the review then the TB should get final say because not trusting anybody's opinion really stalls things.

Nichols says that work needs to be done on the quench protection system for the SS.

Important: Nichols says that the KEY to this job is having the drawings, which is more important than getting permission.

Nichols: Roy said that we should have a CM session to talk about apparatus controls software (engineering?)
Nichols: need RF infrastructure update at CM. Enough design work has been done to identify issues.

Long: we need to discuss the hall layout.

NEW ACTION: Blondel wants Nichols to send the MICO meeting style around to the EB.

Long wants a CM session to review the schedule and 'not blush'. Blondel and Long think this needs to be a TB issue where the session is long enough so that way it can be a "Microsoft Project"-type activity. Regular MICO meetings don't look long enough into the future to address this. Consensus that we shouldn't miss this opportunity to come up with an updated schedule.

NEW ACTION: Nichols to tell CM organizers how much time is needed for schedule discussion

runplan already discussed

NEW ACTION: Blondel and Long to review integration physicist role to determine if it addresses needed tasks.

MICE magnet group

A team of Roy Preece and Mike Courthold which ensures the coordination of magnet issues across collaboration.

FAC papers and MICE MOU

Bonesini prepared MICE MOU. It's being iterated. It's been submitted to John Womersley and somebody else for feedback to check that STFC management was happy with it.

One issue: when we list financial values, it will be obvious to everybody that we aren't talking about same thing. Bonesini talks about spending up to today without provisions for next years.

Confusion over if this is future or past money. We need to consistently include cost of postdocs, students, etc.. Bross asks if this include scientific staff and faculty, which Blondel says would be too hard to assess.

Bonesini suggests to clarify the issue of spent versus globally allocated funds putting appropriate comments in the appendix. Suitable inputs from all parties (INFN, UniGE, US (DOE/NSF), UK, Japan) is needed before the next MICE CM.

ACTION ON KAPLAN: (and Bross) to come up with one number for US for the MOU.

After same comments from Zisman, do we want value or cost?

There is a long discussion on specific costs. Blondel says we can't do this well over phone and need to estimate to the best of our knowledge what the VALUE is if we started from scratch.

SS repair plan status

Already done in TB. Steve Virostek had produced repair plan by Jan. 7, MAP and MZ looked at it. Not ready for full approval. It has gone through the MAP TB. It'll go through another iteration before it gets finalized.

coupling coils

Zisman says that the approach is that Steve Virostek wrote down flowcharts for 5 scenarios ranging from stay the course to startover. The three middle ones are varying degrees of involvement with China. Steve Virostek, Derun Li, and Zisman met about the possibilities. The middle ground is that some stuff is done in China and some in US. Where in the US is a question. The idea is to have crude ideas of when one might be able to deliver and the respective cost. MAP is still the one providing this. Berkeley is only an arm of MAP. The HIT contribution is in doubt. Is there a more robust way to build the CCs? Can this be done at FNAL?

Nichols says the CC need more people and oversight, to which Zisman replies that he's just exploring the realm of possibilities without judgment.

Bross says that Fermilab people say these are difficult magnets to build regardless of who builds them. Experts at FNAL have said it would be hard to build if it were at FNAL, so Bross is worried about the DAQ, scanning voltmeter, etc. since it's 80 km of fiber with a lot of inductance so is hard to test.

Zisman says that QiHuan company has been a shining example in some respects since they've been the quickest to start winding magnets. Bross and Zisman agree they are a good asset to MICE. But on the other hand, HIT which has no progress on the welding and some progress on the testing tasks.

Editorial board

We've asked Bonesini who agreed to head the Editorial board

Plots that can be shown to outside world should get a MICE logo, so people know what to show. We need to avoid contradicting plots at meetings. Palladino to look at MICE note 95 to determine if something needs to be changed about our policy towards new plots. Cobb notes that plots are in flux so it's hard to have forever approved plots.

ACTION: Organize session between Palladino, Cobb and Blondel at the CM.